The interest rate for a savings fund was 3.20% compounded semi-annualy for the first 4 years and then 4.90% compounded semi-annualy for the second 4 years. Harsha invested $2,450 at the end of each 6 months for 8 years.

a. Calculate the accumulated value of her investment after the first 4 years.

b. Calculate the accumulated value of her investment at the end of 8 years.

c. How much interest did she earn in her investment over the 8 years.
Answer #1

Ans:- we will use the FV function of excel to find the future value.

For FV at the end of 4 years, Rate=3.20%/2, Nper=4*2=8, Pmt=-$2450, PV=0

For FV at the end of 8 years, Rate=4.90%/2, Nper=4*2=8, Pmt =-$2,450, PV= -$20,733.43 ( it is accumulated at the end of year 4, so it will become PV at year 4)

G10 A B с 1 N 20733.43 3 Ans:-(a) FV at the end of 4 years 4 5 Ans:-(b) FV at the end of 8 years 46528.89 6 7 8 Ans:-(c) Tota

G10 fac A B C 1 2 3 Ans:-(a) FV at the end of 4 years =FV(3.2%/2,8,-2450,0) 4 5 Ans:-(b) FV at the end of 8 years =FV(4.9%/2,

Note:- Total interest earned over 8 years is calculated by FV at the end of year 8 - Total amount invested

= $46,528.89 - $2,450*16 = $7,328.89.